OpenStudy (anonymous):

how do i get only 1 trig function in: 2cos^-1(y)-sin^-1(y)=0?

zepdrix (zepdrix):

These are the inverse functions? Or did you mean to write \(\Large\rm \frac{1}{\cos y}\)?

OpenStudy (anonymous):

thse are inverse

zepdrix (zepdrix):

Ok ok I think I got it. Boy you've got some tricky questions tonight :D lol

zepdrix (zepdrix):

Let this inverse function be equal to some arbitrary angle, \[\Large\rm \arccos(y)=\theta\]We'll write it in terms of cosine,\[\Large\rm \cos(\theta)=y\]Recall that sine and cosine are `co-functions`, therefore,\[\Large\rm \cos(\theta)=\sin\left(\frac{\pi}{2}-\theta\right)=y\]

zepdrix (zepdrix):

Write it back in terms of the inverse function,\[\Large\rm \arcsin(y)=\frac{\pi}{2}-\theta\]Solving for the angle theta,\[\Large\rm \theta=\frac{\pi}{2}-\arcsin(y)\]And relating this to what we stated earlier,\[\Large\rm \arccos(y)=\theta=\frac{\pi}{2}-\arcsin(y)\]

zepdrix (zepdrix):

So that's how we'll replace our arccos(y) with something involving arcsin(y). Any confusion there? :o

OpenStudy (anonymous):

it seems so simple now. i couldn't figure it out before, but this make sense

OpenStudy (anonymous):

thanks alot zepdrix

zepdrix (zepdrix):

np \c:/ hmm neat problem, i'd never really thought about switching inverses like that before

OpenStudy (alekos):

brilliant!

OpenStudy (anonymous):

its just the 1st part of the problem, cuz its for finding area between curves

zepdrix (zepdrix):

Ooo interesting +_+